Home‎ > ‎Reports‎ > ‎Crew101‎ > ‎

day07


Crew 101 Log Book for March 18, 2011

Photos of the day





Commander's Check-In Report

Tom Caillouet Reporting

Crew Physical Status: Good

Time Departed/Returned from EVA:

EVA 7: 1217 - 1358

EVA 8: 1652 - 1815

Number Of reports:7: (Commander, Journalist, Engineering, 2 EVA, Science, Chef)

Brief description of Field Mission Results: EVA 7 collected partial results for communication ability along the length of Cow Dung Road. EVA 8 collected three biology samples for testing.

Plans for Tomorrow: Reattempt EVA 7 plans for full success. Complete analysis of EVA 8 samples. EVA to improve repeater range by increasing antenna length.



Commander's Report

Tom Caillouet Reporting

It has almost become routine to be out here. No longer do I expect a loud city around me when I wake up, just complete isolation. Along with this routine, we become more efficient in carrying out our work. Prepping for EVAs has become a much more streamlined process as we grow more comfortable with the suits and radio equipment. We rose a bit late today due to a late St. Patrick's Day night celebrating with Irish-themed movies and card games. I had intended to send EVA 7 out at 11 am, but Chrissy and I didn't make it into the airlock until a bit after 11. This caused no output losses on the day, the schedule was merely shifted back an hour.

EVA 7 suffered some setback due to continued power reliability issues with our GPS units. However, even without the antenna at full length, our repeater set up allowed communication, albeit not always very clear, for the entirety of Cow Dung Trail north of the Hab. To gain a reasonable signal to the Hab, the EVA crew had to climb nearby hills about a mile and a half from the Hab, but this is still a great performance and safety gain for the remaining EVA crews. A future EVA will attempt to quantify the radio transmission quality better, so another EVA can test quality with the repeater antenna extended. EVA 8 was accomplished with out any major setbacks. Its goal was to gather soil samples for biological analysis. While the targeted area was not reached, several samples were collected from other stream beds, so the target area will be reached in a later EVA.


Other than EVAs, there have not many other scientific efforts today since all previous samples have been analyzed. In Hab analysis of todays samples will begin tomorrow while we send out EVAs to once again map the northern portion of the trail. If that is successful, and afternoon EVA will boost the repeater antenna length. A separate afternoon EVA collecting geology samples is also possible, but dependent on the rate we get through working with today's samples.


Engineering Report

Cody Hall Reporting

Generator/Electricity:

Kitty  (Diesel Generator) (Used? Yes/No): Yes

Honey  (Gas Generator) (Used? Yes/No): No

Battery Charge Level:

o     V everstart

o     V armor plate

o     V green


Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0.75

Propane Tank (Full/0.75/0.5/0.25/Empty): 0.25

Gasoline Tank (Full/0.75/0.5/0.25/Empty): temperature too low to check after use

Notes/Comments: refilled red hab gasoline container


Oil:

Kitty Oil Quantity (# of Quarts in storage): Leave it to DG

Honey Oil Quantity (# of Quarts in storage): Leave it to DG

ATV Oil Quantity (# of Quarts in storage): 6 quarts on the

engineering work bench

Car Oil Quantity (# of Quarts in storage): Leave it to DG

Notes/Comments:


Water Status:

Outside Potable Water Tank Level (inches from bottom):  18

Trailer Potable Water Tank Level (inches from bottom): 6

Hab Potable Water Tank Level (inches from bottom): 10

Potable Water Meter Reading: 436411


Notes/Comments :


GreenHab:

Condition of plants in Tank 1: unchanged

Condition of plants in Tank 2: unchanged

Condition of plants in Tank 3: water level dropped 0.5 inches

Greenhab notes/ comments:


Observatory:

Telescope Used? (Yes/no) No

Observatory Notes:


Transportation:

Hab Car (NOT RENTAL)

Used (yes/no): No

Oil Checked (yes/no): No

Tire Status: Not checked

Odometer at end of day: 123688.4

Notes/Comments on Hab Car:


ATV 1 (Opportunity):

Used: yes

Oil Checked: no

Fuel Consumed: replaced

Tire Status: ok

Comments:


ATV 2 (Spirit):

Used: yes

Oil Checked: no

Fuel Consumed: replaced

Tire Status: ok

Comments:


ATV 3 (Viking I):

Used: no

Oil Checked: no

Fuel Consumed: none

Tire Status: ok

Comments


ATV 4 (Viking II):

Used: no

Oil Checked: no

Fuel Consumed: none

Tire Status: ok

Comments:

ATV Comments: note that Spirit has been moved to ATV 2 in report


Heating and Ventilation:

Thermostat setting upstairs (Farenheit): 68

Actual temperature upstairs (Farenheit): 71

Thermostat setting downstairs (Farenheit): 55

Actual temperature downstairs (Farenheit): 61

Notes/Comments:


Computers/Networking Infrastructure: working

Notes/comments on internet/computers:


General Engineering

Notes/Comments:



Science Report

Tom Caillouet Reporting

Georgia Tech crews at MDRS have used a HAM radio setup for six previous rotations. Crew 101 continues the use of the system to improve EVA efficiency and to study the range of communication of the equipment in the Mars-like enviroment of MDRS. EVA 7 on March 18, 2011

was a range testing expedition along the main trail next to the Hab.

Comms checks were planned to occur about every three minutes. At points where the voice quality became greatly reduced, GPS coordinates were to be taken at sample spots. Clear main road communications were maintained to about a half mile north of Olympus Mons. Further along the road, Hab transmissions could still reach the crew members on EVA, but crewmembers in the field could not achieve transmissions directly to the Hab. However, by climbing nearby hills, transmissions from the field could be received by the Hab. The first location at which this was accomplished was at 12 N 0518504 4252834. Unfortunately, the GPS lost power after the first reading, but the crew continued to gather rough performance data for the rest of the trail.

As the trail continues, it increases in elevation and eventually surpasses the altitude of the transceiver on Radio Ridge. Two way road communication was never re-established, but climbing nearby hills usually allowed consistent communications, even with large geographic features between the crew and the repeater. The effect of altitude means that in relatively flat areas, EVA crews are incapable of transmitting to the Hab. Hab transmissions to the EVA crew were received with quite reasonable strength and clarity all the way to Cow Dung Reservoir, which the EVA crew estimated to be about 3 miles from the Hab. At this point, the EVA crew was unable to transmit to the Hab, and also unable to reach a hill top to transmit from. The closest proven transmission point was about a quarter mile away.



EVA 7 Report

Tom Caillouet Reporting

EVA Time: 12:17 - 13:58

Mode of Transportation: ATV

EVA Procedure: Crew members will travel on ATVs along Cow Dung Road, stopping about every three minutes for a radio signal strength check.

EVA Results: Traveled to Cow Dung Resevoir. Radio signal testing along main trail only partially completed due to GPS failure.

Lessons Learned: Rechargeable batteries do not provide sufficient charge to run GPS for any reasonable amount of time.



EVA 8 Report

Christina Graves Reporting

EVA Time: 1652-1815

EVA Crewmembers: Cody (KJ4RLF), Christina

Base Communications: Tommy (KJ4KAC)

EVA Procedure: The crew set out to find Cow Dung Reservoir for soil samples for geology and biology experiments, but were unable to locate the site. Soil samples were taken from alternative locations, instead. The sites were chosen based on characteristics of water flow. General site locations were north and north east of the hab, and were predetermined, and based on a satellite image map of the area.  The team was not to deviate from established ATV paths, and ATV helmets were exchanged with bubble helmets prior to sample collection and on-foot exploration of sites.


EVA Results: Soil samples were returned from two sites located at 12S 0518280, 4255774 UTM (Site 1) and 12S 0518175, 4253614 UTM (Site 2).  Accuracy of the first GPS location was 4m, with an altitude of 1370m. A third sample (consisting of a 1 lb. rock) was collected at 12S 0518584, 4252862 UTM, with a GPS accuracy of 5m, and an altitude of 1359.

Lessons Learned: Accurately map and locate destinations prior to ATV; or, take along a crew member who is familiar with the destination of interest.


Chef's Report

Cody Hall Reporting

1a) Today was a non-cooking day.

1b) Was today a special day celebrated at the main meal with special food or activity?

No

2) List the foods served at today's main meal (usually the evening meal), giving full names of each.

Rice with Sweet Pepper Beef Steak

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)

Rice with Sweet Pepper Beef Steak

4) Recipes for anything more complicated than following package directions exactly, or rehydrating.

5) Lessons learned (if any)

Don’t make entire can of food. Crew cannot eat it all.

6) Comments/questions for Kim and Jean, the food study investigators

None


Journalist Report

Christine Redmond Reporting

Last night crew 101 had a blast chatting with the NASA INSPIRE students via live chat! We had fun participating in their “pre-chat poll”, arguing amongst ourselves prior to the chat starting whether fast food and pizza  or family and friends is missed more here on Mars. It was exciting seeing the chat map fill with students from all over the United States and Puerto Rico! It was impressive hearing from the INSPIRE students as they described their experiment proposals to the chat group. After the chat we had a great St. Patty’s day celebration complete with Irish soda bread and Sheppard’s pie! We were all impressed with the quality of the Sheppard’s pie considering it was made up of entirely dehydrated ingredients.

This morning Tommy and I went on an exploratory EVA and tested the range of the radios in the process. I had a blast and became much more comfortable operating the ATVs. Last night one of the INSPIRE students ask the crew the question, “What made you want to participate in MDRS?” My quick and simple response was “because it is fun!” Tommy expanded on this by explaining that many of us are aerospace engineers and many of us will never become astronauts. He views MDRS as an opportunity for us as engineers to further understand the environment and limitations in which we are designing for. Today’s EVA further confirmed Tommy’s statement for me. When Tommy and I stopped to take our first GPS coordinate during our EVA we could not get our GPS to turn on. We tried to turn on our backup GPS which also failed. So we were tasked with the daunting task of changing the GPS batteries while wearing our thick, space gloves. After today’s EVA I can say that I am an aerospace engineer who better understands the usability issues astronauts on Mars would face.

During our exploratory EVA we discovered an area with some very interesting geography and stream beds for future science EVAs. Today was a jam packed day! As soon as we arrived back at the Hab, Cody and Christina were ready to suit up for EVA #8. Cody and Christina traveled to a nearby reservoir to collect biology samples. Now everyone is back at base and ready for dinner! Today is a pre-prepared day so we are all hoping that we will not have a “Mesquite BBQ Rice and Beans with Chicken” repeat!